# Bayesian Experimental Design via Contrastive Diffusions

### v1

@@ 4 EIG optimization through sampling
Iterative sampling operators, as introduced in Marion et al., (2024), are mappings from a space, to be specified below, to a space of probabilities. In our BOED setting, we consider two such operators. The first one is defined, for each ${\boldsymbol{\xi}}$, through a sequence over $s$ of parameterized functions from ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ to ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ and denoted by $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}(p,{\boldsymbol{\xi}})$. Sampling is defined as the outcome in the limit $s\rightarrow\infty$ or for some finite $s=S$ of the following process starting from $p^{(0)}\in{\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ and iterating
$p^{(s+1)}=\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}(p^{(s)},{\boldsymbol{\xi}})\;.$ (15)
This formalism includes normalizing flows (Papamakarios et al.,, 2021) for which $p^{(s)}$ is explicit. More generally, we also consider situations in which $p^{(s)}$ is represented by a random variable ${\mathb{X}}_{s}\sim p^{(s)}$. For example, in density-based BOED, when considering a Langevin diffusion converging to $p_{\boldsymbol{\xi}}$, for a given ${\boldsymbol{\xi}}$, $p^{(s+1)}=\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}(p^{(s)},{\boldsymbol{\xi}})$ is a shorthand for its Euler discretization,
${\mathb{y}}^{(s+1)}$ $={\mathb{y}}^{(s)}-\gamma_{s}\nabla_{\mathb{y}}V({\mathb{y}}^{(s)},{\boldsymbol{\theta}}^{(s)},{\boldsymbol{\xi}})+\sqrt{2\gamma_{s}}{\mathb{B}}_{{\mathb{y}},s}$
${\boldsymbol{\theta}}^{(s+1)}$ $={\boldsymbol{\theta}}^{(s)}-\gamma_{s}\nabla_{\boldsymbol{\theta}}V({\mathb{y}}^{(s)},{\boldsymbol{\theta}}^{(s)},{\boldsymbol{\xi}})+\sqrt{2\gamma_{s}}{\mathb{B}}_{{\boldsymbol{\theta}},s}$
where ${\mathb{B}}_{{\mathb{y}},s}$ and ${\mathb{B}}_{{\boldsymbol{\theta}},s}$ are realizations of independent standard Gaussian variables, $\gamma_{s}$ is a step-size and $V({\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})=-\log p({\boldsymbol{\theta}})-\log p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ is the $p_{\boldsymbol{\xi}}$ potential, $p_{\boldsymbol{\xi}}({\mathb{y}},{\boldsymbol{\theta}})\propto\exp\left(-V({\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})\right)$. The dynamics induced lead to samples from $p_{\boldsymbol{\xi}}$ for $s\rightarrow\infty$. In the following, we will thus use the notation $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}$ to mean that we have access to samples from $p^{(s+1)}$, which is equivalent to apply $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}$ to an empirical version of $p^{(s)}$ built from samples $\{({\mathb{y}}^{(s)}_{i},{\boldsymbol{\theta}}^{(s)}_{i})\}_{i=1:N}$. Similarly for a given posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, we can produce samples $\{{\boldsymbol{\theta}}^{\prime(s+1)}_{j}\}_{j=1:M}$, with a fixed ${\mathb{y}}_{i}$, using the following updating,
${\boldsymbol{\theta}}^{\prime(s+1)}$ $={\boldsymbol{\theta}}^{\prime(s)}-\gamma^{\prime}_{s}\nabla_{\boldsymbol{\theta}}V({\mathb{y}}_{i},{\boldsymbol{\theta}}^{\prime(s)},{\boldsymbol{\xi}})+\sqrt{2\gamma^{\prime}_{s}}{\mathb{B}}_{{\boldsymbol{\theta}}^{\prime},s}\;,$ (16)
while the expected posterior $q_{{\boldsymbol{\xi}},N}$ can be sampled with
@@
For a sampling operator of the expected posterior general form (13), we need to extend the definition in Marion et al., (2024) by adding a dependence on some distribution $\rho\in{\cal P}({\cal Y})$ for the conditioning part. The second sampling operator is defined, for some given ${\boldsymbol{\xi}}$ and $\rho$, through a sequence over $s$ of parameterized functions from ${\cal P}({\boldsymbol{\Theta}})$ to ${\cal P}({\boldsymbol{\Theta}})$ and denoted by $\Sigma_{s}^{{\boldsymbol{\theta}}^{\prime}}(q,{\boldsymbol{\xi}},\rho)$. The sampling operator is defined as the outcome of the following process starting from $q^{(0)}\in{\cal P}({\boldsymbol{\Theta}})$ and iterating
@@
Figure 2: Source localisation example. Prior (left) and expected posterior (right) samples at experiment $k$. Final ${\boldsymbol{\xi}}_{k}^{*}$ (orange cross) at the end of the optimization sequence ${\boldsymbol{\xi}}_{0},\cdot,{\boldsymbol{\xi}}_{T}$ (blue crosses). This optimization "contrasts" the two distributions by making the expected posterior "as different as possible" from the prior.
@@ 5 Single loop contrastive EIG optimization
The perspective of optimization through sampling leads naturally to a nested loop procedure. An inner loop is performed, iterating sampling operators, to reach good approximations of $p_{\boldsymbol{\xi}}$ and $q_{{\boldsymbol{\xi}},N}$ via samples used to evaluate gradient (10). It results a nested loop algorithm in which the design is updated only after completion of the inner loop, which in our BOED context involves two samplers, as specified in Section 4 and summarized in Algorithm 1. More efficient procedures consist of iterating alternately between the sampling operator, e.g. one Langevin step, and the design update, via one iteration of stochastic gradient descent (SGD) or another optimizer. In particular, Marion et al., (2024) point out the link to bi-level optimization when sampling can be cast as an optimization over the space of distributions. For instance, the Langevin diffusion dynamics follow a gradient flow of the KL with respect to the Wasserstein-2 distance (Jordan et al.,, 1998). This point of view suggests to transform the nested loop Algorithm 1 into a single loop summarized in Algorithm 2. At each optimization step over ${\boldsymbol{\xi}}$, the sampling operators are applied only once using the current ${\boldsymbol{\xi}}$, which is updated in turn, etc. Algorithms 1 and 2 are in a general form. In this section, we specify the steps that are further illustrated in Section 6. Regarding the choice of sampling operators, this approach can handle both traditional density-based sampling where an expression of the target distribution is available and data-based sampling where only training samples are available and a generative model is used to generate additional similar samples. Examples of both cases are specified below.
@@
Sampling operators examples for BOED.
@@
Examples of data-based samplers are generative models such as denoising diffusions and others (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020; Song et al.,, 2021) that use score-matching techniques (Hyvärinen,, 2005; Vincent,, 2011). Diffusion models, over a time index $t$, learn a score function $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)})$ using a neural network $s_{\phi}({\boldsymbol{\theta}}^{(t)},t)$ with parameters $\phi$. Given this learned approximation, it is possible to generate new samples of ${\boldsymbol{\theta}}$ by running a backward diffusion process using $s_{\phi}({\boldsymbol{\theta}}^{(t)},t)$ instead of the true score. In BOED, we are interested in sampling from a conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ with the following objectives. First we need to sample from the expected posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}})$ which requires conditioning on the observation ${\mathb{y}}$. Second, in sequential design problems (see section 6.1 and Appendix D), we need to condition on the history of observations ${\mathb{D}}_{k-1}$ and produce samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$. Both these issues can be tackled in the framework of denoising diffusions with linear inverse problems (Daras et al.,, 2024). When the likelihood corresponds to a measurement ${\mathb{Y}}$ with ${\mathb{Y}}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\boldsymbol{\eta}}$ and ${\boldsymbol{\eta}}\!\sim\!\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}})$, inspiring recent attempts, such as (Corenflos et al.,, 2024; Cardoso et al.,, 2024), have addressed the problem of sampling efficiently from $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ using only the learned score $s_{\phi}({\boldsymbol{\theta}}^{(t)},t)$, without the need for any kind of retraining. Denoising diffusions include several variants but some formulations consist of running a stochastic differential equation (SDE), see Appendix C. For conditional sampling, this would mean running an SDE with a conditional score $\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$, which is intractable. See (42) and more details in Appendix C.2. As a solution, Dou and Song, (2024) propose a method named FPS that approximates $p({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$ by $p({\boldsymbol{\theta}}^{(t)}|{\mathb{y}}^{(t)},{\boldsymbol{\xi}})$ with ${\mathb{y}}^{(t)}$ the noised observation at time $t$. As the score $\nabla_{\boldsymbol{\theta}}\!\log p({\boldsymbol{\theta}}^{(t)}|{\mathb{y}}^{(t)},{\boldsymbol{\xi}})$ can be written as $\nabla_{\boldsymbol{\theta}}\!\log p({\boldsymbol{\theta}}^{(t)})\!+\!\nabla_{\boldsymbol{\theta}}\!\log p({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$, we can leverage the learned score $s_{\phi}({\boldsymbol{\theta}}^{(t)},t)$ and the closed form of $\nabla_{\boldsymbol{\theta}}\log p({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$ to sample approximately from $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ using a backward SDE with the approximate score, see (44) in Appendix C.2. This allows to sample efficiently from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and, using $\nabla_{\boldsymbol{\theta}}\log q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})=\sum_{i=1}^{N}\nu_{i}\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, from the expected posterior with the extension of (44) below, where ${\mathb{y}}_{i}^{(t)}$ is the noised ${\mathb{y}}_{i}$ at time $t$,
$d{\boldsymbol{\theta}}^{{}^{\prime}(t)}=\left[-\frac{\beta(t)}{2}{\boldsymbol{\theta}}^{{}^{\prime}(t)}-\beta(t)\sum_{i=1}^{N}\nu_{i}\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{{}^{\prime}(t)}|{\mathb{y}}_{i}^{(t)},{\boldsymbol{\xi}})\right]dt+\sqrt{\beta(t)}d{\mathb{B}}_{t}\;.$ (19)
An additional resampling SMC-like step can also be added as explained in Appendix E. The approach allows to handle new sequential data-based BOED tasks as illustrated in Section 6.3.
@@
Optimizing ${\boldsymbol{\xi}}$ using the gradient expression (10) encourages to select a ${\boldsymbol{\xi}}$ that gives either high probability $p({\mathb{y}}_{i}|{\boldsymbol{\theta}}_{i},{\boldsymbol{\xi}})$ to samples $({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})$ from $p_{\boldsymbol{\xi}}$ or low probability $p({\mathb{y}}_{j}|{\boldsymbol{\theta}}^{\prime}_{j},{\boldsymbol{\xi}})$ to samples ${\boldsymbol{\theta}}^{\prime}_{j}$ from $q_{{\boldsymbol{\xi}},N}$. This contrastive behaviour is also visible in (2) where the EIG is defined as the mean over the experiment outcomes of the KL between posterior and prior distributions. The expected posterior $q_{{\boldsymbol{\xi}},N}$ is then used as a proxy to the intractable posterior, to perform this contrastive optimization. Figure 2 provides a visualization of this contrastive behavior in the source localization example of Section 6.2. It corresponds to set the next design ${\boldsymbol{\xi}}$ to a value that eliminates the most parameter ${\boldsymbol{\theta}}$ values (right plot) among the possible ones a priori (left plot). This is analogous to Noise Constrastive Estimation (Gutmann and Hyvärinen,, 2010) methods where model parameters are computed so that the data samples are as different as possible from the noise samples. Additional illustrations are given in Appendix Figure 5.

### v2

@@ 4 EIG optimization through sampling
Iterative sampling operators, as introduced in Marion et al., (2025), are mappings to a space of probabilities. In our BOED setting, we consider two such operators. The first one is defined, for each ${\boldsymbol{\xi}}$, through a sequence over $s$ of functions from ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ to ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ and denoted by $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}(p,{\boldsymbol{\xi}})$. Sampling is defined as the outcome in the limit $s\rightarrow\infty$ or for some finite $s=S$ of the following process starting from $p^{(0)}\in{\cal P}({\boldsymbol{\Theta}}\times{\cal Y})$ and iterating
$p^{(s+1)}=\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}(p^{(s)},{\boldsymbol{\xi}})\;,$ (15)
where $p^{(s)}$ can be explicit, e.g. $p^{(s)}$ is a Gaussian distribution with parameters depending on $s$, or represented by a random variable ${\mathb{X}}_{s}\sim p^{(s)}$. For example, in density-based BOED, for ${\mathb{X}}_{s}=({\mathb{Y}}_{s},{\boldsymbol{\theta}}_{s})$, we consider the Euler discretization of the Langevin diffusion converging to $p_{\boldsymbol{\xi}}$
${\mathb{y}}^{(s+1)}$ $={\mathb{y}}^{(s)}-\gamma_{s}\nabla_{\mathb{y}}V({\mathb{y}}^{(s)},{\boldsymbol{\theta}}^{(s)},{\boldsymbol{\xi}})+\sqrt{2\gamma_{s}}{\mathb{B}}_{{\mathb{y}},s},$ (16)
${\boldsymbol{\theta}}^{(s+1)}$ $={\boldsymbol{\theta}}^{(s)}-\gamma_{s}\nabla_{\boldsymbol{\theta}}V({\mathb{y}}^{(s)},{\boldsymbol{\theta}}^{(s)},{\boldsymbol{\xi}})+\sqrt{2\gamma_{s}}{\mathb{B}}_{{\boldsymbol{\theta}},s}.$
where ${\mathb{B}}_{{\mathb{y}},s}$ and ${\mathb{B}}_{{\boldsymbol{\theta}},s}$ are realizations of independent standard Gaussian variables, $\gamma_{s}$ is a step-size and $V({\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})=-\log p({\boldsymbol{\theta}})-\log p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ is the $p_{\boldsymbol{\xi}}$ potential, $p_{\boldsymbol{\xi}}({\mathb{y}},{\boldsymbol{\theta}})\propto\exp\left(-V({\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})\right)$. The dynamics induced lead to samples from $p_{\boldsymbol{\xi}}$ for $s\rightarrow\infty$. In the following, we will thus use the notation $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}$ to mean that we have access to samples from $p^{(s+1)}$, which is equivalent to apply $\Sigma_{s}^{{\mathb{Y}},{\boldsymbol{\theta}}}$ to an empirical version of $p^{(s)}$ built from samples $\{({\mathb{y}}^{(s)}_{i},{\boldsymbol{\theta}}^{(s)}_{i})\}_{i=1:N}$. Similarly, we can produce samples $\{{\boldsymbol{\theta}}^{\prime(s+1)}_{j}\}_{j=1:M}$ from the pooled posterior $q_{{\boldsymbol{\xi}},N}$, using its score expression, via the updating,
@@
For a sampling operator of the pooled posterior general form (13), we need to extend the definition in Marion et al., (2025) by adding a dependence on some distribution $\rho\in{\cal P}({\cal Y})$ for the conditioning part. The second sampling operator is defined, for some given ${\boldsymbol{\xi}}$ and $\rho$, through a sequence over $s$ of parameterized functions from ${\cal P}({\boldsymbol{\Theta}})$ to ${\cal P}({\boldsymbol{\Theta}})$ and denoted by $\Sigma_{s}^{{\boldsymbol{\theta}}^{\prime}}(q,{\boldsymbol{\xi}},\rho)$. The sampling operator is defined as the outcome of the following process starting from $q^{(0)}\in{\cal P}({\boldsymbol{\Theta}})$ and iterating
@@
Figure 2: Source localisation example. Prior (left) and pooled posterior (right) samples at experiment $k$. Final ${\boldsymbol{\xi}}_{k}^{*}$ (orange cross) at the end of the optimization sequence ${\boldsymbol{\xi}}_{0},\cdot,{\boldsymbol{\xi}}_{T}$ (blue crosses). This optimization "contrasts" the two distributions by making the pooled posterior "as different as possible" from the prior.
@@ 5 Single loop contrastive EIG optimization
The perspective of optimization through sampling leads naturally to a nested loop procedure. An inner loop is performed to reach good approximations of $p_{\boldsymbol{\xi}}$ and $q_{{\boldsymbol{\xi}},N}$ using two samplers as specified in Section 4 and summarized in the nested loop Algorithm 1. Considering sampling as an optimization over the space of distributions (Marion et al.,, 2025), a more efficient single loop procedure can be derived. As illustrated in the single loop Algorithm 2, at each optimization step over ${\boldsymbol{\xi}}$, the sampling operators are applied only once using the current ${\boldsymbol{\xi}}$, which is updated in turn, etc. Sampling operators can be derived from traditional density-based sampling, like in (16) and (17), where an expression of the target distribution is required to compute the score, and also from data-based sampling where only training samples are available. In the latter case, conditional score-based generative models have emerged as a very active field of research. We explicit below how a recent such framework proposed by Dou and Song, (2024) can be used in our setting.
Data-based samplers.
In BOED, we are interested in sampling from a conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ with the following objectives. First we need to sample from the pooled posterior $q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}})$ which requires conditioning on the observation ${\mathb{y}}$. Second, in sequential design problems (see section 6.1 and Appendix D), we need to condition on the history of observations ${\mathb{D}}_{k-1}$ and produce samples from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$. Both these issues can be tackled in the framework of diffusion models for inverse problems (Daras et al.,, 2024). When the likelihood corresponds to a linear measurement ${\mathb{Y}}$ with ${\mathb{Y}}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\boldsymbol{\eta}}$ and ${\boldsymbol{\eta}}\!\sim\!\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}})$, inspiring recent attempts, such as (Corenflos et al.,, 2025; Cardoso et al.,, 2024), have addressed the problem of sampling efficiently from $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ using only the pre-trained score $s_{\phi}({\boldsymbol{\theta}},t)$ of a diffusion model, without the need for any kind of retraining (see Appendix C for details). For conditional sampling, this would mean running an SDE with a conditional score $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$, which is intractable. See (43) and Appendix C.2. As a solution, Dou and Song, (2024) propose a method named FPS that approximates $p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$ by $p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}}^{(t)},{\boldsymbol{\xi}})$ with ${\mathb{y}}^{(t)}$ the noised observation at time $t$. As the score $\nabla_{\boldsymbol{\theta}}\!\log p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}}^{(t)},{\boldsymbol{\xi}})$ can be written as $\nabla_{\boldsymbol{\theta}}\!\log p_{t}({\boldsymbol{\theta}}^{(t)})\!+\!\nabla_{\boldsymbol{\theta}}\!\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$, we can leverage the learned score $s_{\phi}({\boldsymbol{\theta}}^{(t)},t)$ and the closed form of $\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$ to sample approximately from $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ using a backward SDE with the approximate score, see (C.2) in Appendix C.2. This allows to sample efficiently from $p({\boldsymbol{\theta}}|{\mathb{D}}_{k-1})$ and, using $\nabla_{\boldsymbol{\theta}}\log q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}}^{\prime})\!=\!\sum_{i=1}^{N}\!\nu_{i}\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, from the pooled posterior with the extension of (C.2) below, where ${\mathb{y}}_{i}^{(t)}$ is the noised ${\mathb{y}}_{i}$ at time $t$ of the forward SDE,
$d{\boldsymbol{\theta}}^{{}^{\prime}(t)}=\left[-\frac{\beta(t)}{2}{\boldsymbol{\theta}}^{{}^{\prime}(t)}-\beta(t)\sum_{i=1}^{N}\nu_{i}\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{{}^{\prime}(t)}|{\mathb{y}}_{i}^{(t)},{\boldsymbol{\xi}})\right]dt+\sqrt{\beta(t)}d{\mathb{B}}_{t}\;,$ (19)
where $t$ above is now flowing backwards from infinity to $t\!\!=\!\!0$. In practice, (19) is solved approximately using a numerical discretization and an initialization of the process with ${\boldsymbol{\theta}}^{{}^{\prime}(T)}\sim{\cal N}(\mathbf{0},{\mathb{I}})$ for some large finite $T$. An additional resampling SMC-like step can also be added as explained in Appendix E. The approach allows to handle new sequential data-based BOED tasks as illustrated in Section 6.3.
@@
Measure 1 2 3 4 5 6
CoDiff .227 .338 .528 .673 .789 .826
Random .168 .275 .350 .391 .421 .463
Table 1: CoDiff and random reconstruction quality comparison with SSIM, in [-1,1], the higher the better.
Density-based samplers.
@@
Optimizing ${\boldsymbol{\xi}}$ using the gradient expression (10) encourages to select a ${\boldsymbol{\xi}}$ that gives either high probability $p({\mathb{y}}_{i}|{\boldsymbol{\theta}}_{i},{\boldsymbol{\xi}})$ to samples $({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})$ from $p_{\boldsymbol{\xi}}$ or low probability $p({\mathb{y}}_{j}|{\boldsymbol{\theta}}^{\prime}_{j},{\boldsymbol{\xi}})$ to samples ${\boldsymbol{\theta}}^{\prime}_{j}$ from $q_{{\boldsymbol{\xi}},N}$. This contrastive behaviour is also visible in (2) where the EIG is defined as the mean over the experiment outcomes of the KL between posterior and prior distributions. The pooled posterior $q_{{\boldsymbol{\xi}},N}$ is then used as a proxy to the intractable posterior, to perform this contrastive optimization. Figure 2 provides a visualization of this contrastive behavior in the source localization example of Section 6.2. It corresponds to set the next design ${\boldsymbol{\xi}}$ to a value that eliminates the most parameter ${\boldsymbol{\theta}}$ values (right plot) among the possible ones a priori (left plot). This is analogous to Noise Constrastive Estimation (Gutmann and Hyvärinen,, 2010) methods where model parameters are computed so that the data samples are as different as possible from the noise samples. Additional illustrations are given in Appendix Figure 6.